# BiSSL: Enhancing the Alignment Between Self-Supervised Pretraining and Downstream Fine-Tuning via Bilevel Optimization

### v4

@@ 4.3 Downstream Task Performance
To assess if the advantages of BiSSL extend beyond classification tasks, we further assess its performance in the context of object detection. Similarly to the setup of [2, 3], we use a Faster-RCNN with a ResNet-50 C4 backbone [67] on the VOC07+12 [58] dataset and report the standard $\text{AP}_{50}$ metric. We refer to Section B.6 in Appendix B for more implementation details. The results in Table 2 support the previous conclusions, indicating BiSSL significantly improves downstream performance.
Figure 2: Top-1 test classification accuracies on the flowers dataset for separate models pretrained with SimCLR for different durations, comparing the conventional SSL and BiSSL training pipelines. BiSSL consistently achieves significantly higher accuracies than the baseline when the pretraining duration is sufficiently high.
Figure 3: Feature visualizations of BiSSL-trained (upper) vs pretext-only (lower) backbones on the flowers dataset. Colors denote respective classes. Details in Section C.8.
Appendix C provides additional benchmark results. Section C.4 shows that BiSSL also outperforms conventional fine-tuning when using a Masked Autoencoder [7] for self-supervised pretraining. Section C.5 extends comparisons to more baselines, with BiSSL again performing best. Lastly section C.7 presents ImageNet fine-tuning results, where BiSSL performs comparably to conventional fine-tuning, which is expected since pretraining is also conducted on ImageNet, leading to strong alignment between the pretext and downstream distributions.
@@ 4.3.1 Varying the pretraining duration
To further assess the robustness of BiSSL, we vary the duration of self-supervised pretraining. Due to computational resource constraints, we adopt a smaller-scale version of the SimCLR setup, using a ResNet-18 backbone and the unlabeled partition of the smaller-scale STL10 [68] dataset for pretraining. Reusing the parameters from Section 4.2.2, BiSSL training corresponds to $100$ conventional pretext epochs using the STL10 dataset. To ensure fair comparison, the pretext-only baselines are accordingly pretrained for $100$ additional epochs. The rest of the setup remains unchanged. We evaluate on the flowers dataset, where BiSSL previously showed notable gains. Figure 3 depicts the final top-1 test accuracies achieved by separate models pretrained for varying durations, with Figure 4 in Appendix C showing the top-5 test accuracies. The results show that BiSSL consistently outperforms the baseline once sufficient pretraining is reached, aligning with the remarks in Section 3.4, suggesting that its advantage stems not from longer training, but rather from the enhanced information sharing it facilitates between the pretext and downstream tasks.
Appendix C presents additional investigative experiments. Section C.1 reports training durations for pretraining and BiSSL, showing that BiSSL requires only a fraction of the computational cost of pretext pretraining. Section C.3 confirms the effectiveness of the BiSSL setup, showing that setting $N_{U}=1$ or excluding the first term of (2) both degrade downstream performance. Section C.6 demonstrates that BiSSL consistently retains is advantage across varying downstream dataset sizes.
@@ 4.4 Visual Inspection of Latent Features
To better assess whether BiSSL nudges the latent features toward being more semantically meaningful for downstream tasks, we qualitatively inspect latent spaces using the t-Distributed Stochastic Neighbor Embedding (t-SNE) [69] method. Specifically, we compare features from SimCLR-pretrained backbones to those derived from BiSSL-trained lower-level backbones, as described in Sections 4.2.1 and 4.2.2 respectively. Figure 3 illustrates the results on the flowers dataset, indicating that BiSSL improves downstream feature alignment. Additional plots on a selection of downstream datasets in Section C.8 of Appendix C reinforce this finding, also demonstrating that this trend consistently persists even for datasets where BiSSL did not impose any classification accuracy improvements. Further details regarding the experimental setup are outlined in Section C.8.
@@ 5 Conclusion
This study introduces BiSSL, a novel training framework that integrates the self-supervised pretext and downstream fine-tuning objectives into a unified bilevel optimization problem. By explicitly modeling the interdependence between pretraining and fine-tuning, BiSSL facilitates enhanced information sharing between these training stages, yielding model initializations that are better aligned with downstream tasks. We present a task-agnostic training algorithm and pipeline that incorporates BiSSL as an intermediate training stage between the self-supervised pretraining and fine-tuning stages. Experiments across multiple pretext and downstream tasks demonstrate that BiSSL significantly improves downstream performance over the conventional self-supervised learning pipeline. Additional exploratory experiments further support the conclusion that BiSSL enhances downstream task alignment. Overall, BiSSL represents a promising step toward training algorithm designs that better align the self-supervised pretraining and downstream fine-tuning stages via bilevel optimization.
Acknowledgments and Disclosure of Funding
@@
The authors would like to thank Sijia Liu and Yihua Zhang (Michigan State University) providing valuable feedback in a discussion, which helped to refine and solidify our perspective on the topic of integrating bilevel optimization in deep learning.
@@ Appendix A Theoretical Insights and Framework Comparisons in BiSSL
A.1 On the Necessity of the Second Term in the Upper-Level Objective
While the inclusion of the first term $\mathcal{L}^{D}({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\phi}_{D})$ in the upper-level objective in (2) is clearly justified, as it couples the two levels together and enables the upper-level to incorporate information from the lower-level training objective, the second term $\mathcal{L}^{D}({\bf\it\theta}_{D},{\bf\it\phi}_{D})$ may at first glance appear redundant. However, this term is crucial for ensuring that the lower-level problem in (3) is non-redundant when using a mean squared difference regularization objective $r({\bf\it\theta}_{D},{\bf\it\theta}_{P})=\frac{1}{2}\norm{{\bf\it\theta}_{D}-{\bf\it\theta}_{P}}_{2}^{2}$, which is the primary regularization objective in our experiments. We here consider the bilevel optimization problem using the aforementioned regularization objective, while also omitting the second term of (2). To simplify notation and focus on the central idea, we omit head parameters in the following derivation, which does not affect the core reasoning.
Let the conventional fine-tuning problem be defined as:
$\min_{{\bf\it\theta}_{D}}\mathcal{L}^{D}({\bf\it\theta}_{D}),$ (8)
and consider the following bilevel optimization problem:
$\underset{\boldsymbol{\theta}_{D}}{\min}\,\,\mathcal{L}^{D}\left(\boldsymbol{\theta}_{P}^{\ast}(\boldsymbol{\theta}_{D})\right)\quad\text{ s.t.}\quad\boldsymbol{\theta}_{P}^{\ast}\left(\boldsymbol{\theta}_{D}\right)\in\underset{\boldsymbol{\theta}_{P}}{\text{argmin}}\,\mathcal{L}^{P}\left(\boldsymbol{\theta}_{P}\right)+\frac{\lambda}{2}\norm{{\bf\it\theta}_{D}-{\bf\it\theta}_{P}}_{2}^{2}.$ (9)
We then have the following result.
Theorem 1: Let $\bar{{\bf\it\theta}}$ be a stationary point of the conventional fine-tuning problem (8). Define:
${\bf\it\theta}_{P}^{\ast}:=\bar{{\bf\it\theta}},\quad{\bf\it\theta}_{D}^{\ast}:=\bar{{\bf\it\theta}}-\frac{1}{\lambda}\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{{\bf\it\theta}=\bar{{\bf\it\theta}}}.$
Then the pair $({\bf\it\theta}_{D}^{\ast},{\bf\it\theta}_{P}^{\ast})$ satisfies the stationary conditions of the bilevel problem (9).
Theorem 1 shows that under the simplified bilevel formulation in Equation (9), any solution $\bar{{\bf\it\theta}}$ to the conventional fine-tuning problem of (8) are recovered by the bilevel optimization in (9). In other words, the bilevel formulation in (9) effectively reduces to the single-level fine-tuning problem.
The flaw of this setup is that the upper-level is solely tasked with influencing the lower-level solution ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$ via ${\bf\it\theta}_{D}$, without any direct incentive to optimize ${\bf\it\theta}_{D}$ itself. Since ${\bf\it\theta}_{D}$ is only directly present in the lower-level, its sole role becomes one of steering the lower-level dynamics such that ${\bf\it\theta}_{P}$ minimizes the upper-level objective, which undermines the point of the bilevel structure.
Including the term $\mathcal{L}^{D}({\bf\it\theta}_{D})$ in the upper-level objective mitigates this issue by enforcing that ${\bf\it\theta}_{D}$ must itself contribute to satisfying the downstream objective, thereby preventing the lower-level solution from being solely steered to fit the upper-level goal. This yields a genuinely bilevel optimization structure, where both ${\bf\it\theta}_{D}$ and ${\bf\it\theta}_{P}$ are subject to interdependent constraints, promoting a more meaningful interplay between the pretext and downstream tasks.
Proof (Theorem 1): By assumption, $\bar{{\bf\it\theta}}$ is a stationary point of the fine-tuning objective in (8), hence it is straightforward to see for the upper-level in (9) that
$\nabla_{{\bf\it\theta}}\mathcal{L}^{D}({\bf\it\theta})|_{{\bf\it\theta}={\bf\it\theta}_{P}^{\ast}}=\nabla_{{\bf\it\theta}}\mathcal{L}^{D}({\bf\it\theta})|_{{\bf\it\theta}=\bar{{\bf\it\theta}}}=\mathbf{0}.$ (10)
Now, defining the lower-level objective as
$G({\bf\it\theta}_{D},{\bf\it\theta}_{P})=\mathcal{L}^{P}\left(\boldsymbol{\theta}_{P}\right)+\frac{\lambda}{2}\norm{{\bf\it\theta}_{D}-{\bf\it\theta}_{P}}_{2}^{2},$ (11)
we similarly evaluate
$\nabla_{{\bf\it\theta}}G({\bf\it\theta}_{D}^{\ast},{\bf\it\theta})|_{{\bf\it\theta}={\bf\it\theta}_{P}^{\ast}}$ $=\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{{\bf\it\theta}={\bf\it\theta}_{P}^{\ast}}+\lambda({\bf\it\theta}_{D}^{\ast}-{\bf\it\theta}_{P}^{\ast})$ (12)
$=\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{\theta=\bar{{\bf\it\theta}}}+\lambda(\bar{{\bf\it\theta}}-\frac{1}{\lambda}\nabla_{{\bf\it\theta}}\mathcal{L}^{P}({\bf\it\theta})|_{\theta=\bar{{\bf\it\theta}}}-\bar{{\bf\it\theta}})$ (13)
$=\mathbf{0}.$ (14)
Hence, the pair $({\bf\it\theta}_{D}^{\ast},{\bf\it\theta}_{P}^{\ast})$ satisfies the stationary conditions of the bilevel problem in (9). ∎
A.2 Derivation of the Implicit Jacobian
@@
$G({\bf\it\theta}_{D},{\bf\it\theta}_{P}):=\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta}_{D},{\bf\it\theta}_{P}).$ (15)
Recalling that $r$ is a convex regularization objective, adequate scaling of $\lambda$ effectively “convexifies” the lower-level objective $G$, a strategy also employed on the lower-level objective in previous works [30, 27, 26]. This is advantageous because assuming convexity of $G$ ensures that for any ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$, there exists a corresponding $\hat{{\bf\it\theta}}_{P}\in\mathbb{R}^{L}$ that satisfies the stationary condition $\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},\hat{{\bf\it\theta}}_{P})=\mathbf{0}$. In other words, we are assured that a potential minimizer of $G({\bf\it\theta}_{D},\cdot)$ exists for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. Now, further assume that the gradient $\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},{\bf\it\theta}_{P})$ is continuously differentiable and that the Hessian matrix $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},\hat{{\bf\it\theta}}_{P})$ is invertible for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. Under these conditions, the implicit function theorem [70, 71] guarantees the existence of an implicit unique and differentiable function ${\bf\it\theta}_{P}^{\ast}:\mathcal{N}({\bf\it\theta}_{D})\to\mathbb{R}^{L}$, with $\mathcal{N}({\bf\it\theta}_{D})$ being a neighborhood of ${\bf\it\theta}_{D}$, that satisfies ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})=\hat{{\bf\it\theta}}_{P}$ and $\nabla_{{\bf\it\theta}_{P}}G(\tilde{{\bf\it\theta}}_{D},{\bf\it\theta}_{P}^{\ast}(\tilde{{\bf\it\theta}}_{D}))=\mathbf{0}$ for all $\tilde{{\bf\it\theta}}_{D}\in\mathcal{N}({\bf\it\theta}_{D})$.
@@
is valid for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. By applying the chain rule, the expression becomes
@@
Recalling that $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))$ is assumed to be invertible, the implicit jacobian (IJ) $\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ can be isolated
@@
$\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ $=-\lambda\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}\quantity(\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+\lambda r\quantity({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})))]^{-1}$
$=-\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}\quantity(\frac{1}{\lambda}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+r\quantity({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})))]^{-1}.$ (16)

### v5

@@ 4.3.1 Object Detection and Semantic Segmentation
We further assess performance in the context of object detection and semantic segmentation. Similar to the setup of Bardes et al. (2022); He et al. (2020), we use for object detection a Faster-RCNN style architecture with a ResNet-50 C4 backbone (Ren et al., 2015) on the VOC07+12 (Everingham et al., 2010) dataset and report the standard $\text{AP}_{50}$ metric. For semantic segmentation tasks, we follow the setup of He et al. (2020); Grill et al. (2020) and use a FCN-16 (Long et al., 2015b) based structure on top of the pretrained ResNet-50 backbone and evaluate on the VOC07+12 (Everingham et al., 2010) and CityScapes (Cordts et al., 2016) datasets, reporting the mean intersection over union (mIoU) scores. We refer to Section B.7 for complete implementation details. The results in Table 2 support the previous conclusions, indicating that BiSSL overall significantly improves downstream performance.
4.4 Extended Evaluation
In this section, we present a series of experiments to examine how the performance of BiSSL compares with other baselines, and to explore how well BiSSL adapts to different configurations and experimental settings.
4.4.1 Comparative Baseline Evaluations
We assess BiSSL relative to several other baseline methods in this section. Due to computational constraints, we restrict evaluation to five downstream datasets: Pets, DTD, VOC07, Flowers, and CUB200. These cover both coarse- and fine-grained classification tasks across diverse and mostly disjoint visual domains. All baselines involving self-supervised pretraining use SimCLR as the pretext task, as described in Section 4.2.1. We briefly summarize the conceptual design of each baseline and defer full implementation details to Section B.9.
Table 3: Comparison of downstream classification accuracies across baseline methods detailed in Section B.9. For each dataset, the highest average accuracy is marked in bold, and the second-highest is underlined.
Pets DTD VOC07 Flowers CUB200
Random Initialization (B.9.1) $53.9\pm 0.8$ $35.6\pm 0.5$ $43.7\pm 0.3$ $43.7\pm 0.5$ $25.7\pm 0.4$
Linear Probing (B.9.1) $58.4\pm 0.2$ $56.3\pm 0.4$ $63.4\pm 0.1$ $64.5\pm 0.1$ $18.5\pm 0.1$
FT $73.5\pm 0.7$ $60.4\pm 0.9$ $71.0\pm 0.1$ $82.6\pm 0.3$ $49.6\pm 0.3$
FT (Zero Head Init) (B.9.1) $75.6\pm 0.6$ $59.9\pm 0.5$ $70.5\pm 0.2$ $82.2\pm 0.6$ $48.4\pm 1.1$
NoisyTune (B.9.2) $73.4\pm 0.3$ $60.0\pm 0.7$ $70.3\pm 0.2$ $82.6\pm 0.2$ $48.9\pm 0.5$
Pretraining on Data Mix (B.9.3) + FT $74.6\pm 0.6$ $60.1\pm 0.6$ $69.6\pm 0.2$ $81.6\pm 0.4$ $50.9\pm 0.7$
Pretext-Downstream Sum (B.9.4) + FT $\underline{76.4\pm 0.4}$ $62.2\pm 0.3$ $67.1\pm 0.1$ $\underline{83.8\pm 0.1}$ $52.9\pm 0.5$
BiSSL (IJ Discarded) (B.9.5) + FT $76.2\pm 0.3$ $\underline{63.0\pm 0.3}$ $\mathbf{71.7\pm 0.1}$ $81.5\pm 0.3$ $\underline{53.1\pm 0.3}$
BiSSL + FT $\mathbf{78.3\pm 0.2}$ $\mathbf{64.2\pm 0.4}$ $\underline{71.5\pm 0.1}$ $\mathbf{84.1\pm 0.2}$ $\mathbf{55.7\pm 0.2}$
We include baselines consisting of a randomly initialized backbone, linear probing, and fine-tuning with a zero-initialized head, all described in Section B.9.1. Then, we further compare BiSSL to methods that similarly introduce an intermediate stage between pretraining and fine-tuning. These include NoisyTune (Wu et al., 2022), which perturbs backbone parameters with noise (details in Section B.9.2), continued pretraining on a mixture of downstream and pretext data (detailed in Section B.9.3), and joint minimization of a weighted sum of the downstream and pretext objectives (detailed in B.9.4). In addition, we also evaluate a BiSSL variant with the first term of the upper-level objective (2) discarded, eliminating the gradient coupling from the lower to upper-level, which effectively reduces to an alternating optimization scheme (details in B.9.5). This variant can itself also be considered an ablation, as it tests whether the increased information sharing imposed by the bilevel structure provides any tangible benefit for the downstream tasks.
Table 3 reports the results for all mentioned baselines, along with the FT and BiSSL+FT results from Table 1 for ease of comparison. As expected, pretrained backbones consistently outperform initializing them randomly, and fine-tuning yields further gains over linear probing. Notably, DTD and VOC07 exhibit the smallest relative drop in performance under linear probing compared to fine-tuning, suggesting that their tasks benefit more directly from the general-purpose representations captured during SSL pretraining. This could stem from the fact that VOC’s labels are relatively coarse-grained, while DTD relies on general texture patterns rather than fine-grained distinctions, in contrast to the fine-grained Pets, Flowers, and CUB200 datasets. BiSSL overall achieves the strongest results, ranking first on four of the five datasets and second on the remaining one. In contrast, NoisyTune fails to improve downstream accuracy in this setting, and the remaining baselines provide gains over conventional fine-tuning less consistently, and when improvements do occur, they are generally smaller than those achieved by BiSSL.
Table 4: Downstream classification performance when using a ViT-backbone with MAE as the pretext task. BiSSL improves performance with statistical significance and reduces variance across all five datasets.
MAE Pets DTD VOC07 Flowers CUB200
FT $81.3\pm 0.5$ $61.3\pm 0.7$ $69.5\pm 0.3$ $81.1\pm 1.1$ $59.0\pm 0.7$
BiSSL+FT $\mathbf{82.9\pm 0.1}$ $\mathbf{62.9\pm 0.4}$ $\mathbf{72.0\pm 0.2}$ $\mathbf{87.3\pm 0.3}$ $\mathbf{65.1\pm 0.4}$
Avg Diff $\mathbf{+1.6}$ $\mathbf{+1.6}$ $\mathbf{+2.5}$ $\mathbf{+6.2}$ $\mathbf{+6.1}$
4.4.2 Transformer-Based Backbone Architecture
Our main experiments establish the effectiveness of BiSSL using ResNet-based architectures. To evaluate whether these benefits extend to other architectures, we test BiSSL with the more recent ViT backbone (Dosovitskiy et al., 2021), trained using the Masked Autoencoder (MAE) pretext task (He et al., 2022). Experimental details are provided in Section B.8. The evaluation uses the same five downstream datasets as in Section 4.4.1. As shown in Table 4, BiSSL delivers consistent and statistically significant improvements across all datasets and reduces variance in downstream accuracy compared to standard fine-tuning.
@@ 4.4.3 Varying the Pretraining Duration
To further assess the robustness of BiSSL, we vary the duration of self-supervised pretraining. Due to computational resource constraints, we adopt a smaller-scale version of the SimCLR setup, using a ResNet-18 backbone and the unlabeled partition of the smaller-scale STL10 (Coates et al., 2011) dataset for pretraining. Reusing the parameters from Section 4.2.2, BiSSL training corresponds to $100$ conventional pretext epochs using the STL10 dataset. To ensure fair comparison, the pretext-only baselines are accordingly pretrained for $100$ additional epochs. The rest of the setup remains identical to as described in Section 4.2. We evaluate on the flowers dataset, where BiSSL previously showed significant gains. Figure 3 depicts the final top-1 test accuracies achieved by separate models pretrained for varying durations. BiSSL consistently outperforms the baseline once sufficient pretraining is reached, consistent with the remarks in Section 3.4.
Figure 2: Top-1 test accuracies on the flowers dataset for models pretrained with SimCLR across different training durations. BiSSL consistently achieves significantly higher accuracies when the pretraining duration is sufficiently high.
Figure 3: Parameter-wise and feature-wise cosine similarities between SimCLR-pretrained and BiSSL-trained backbones with their respective fine-tuned counterparts. Results are averaged over the 12 downstream classification datasets listed in Table 1, and values for each individual dataset are listed in Table 9 of Section B.10.1. The results indicate that BiSSL increases downstream alignment of backbones.
Additional Results
We provide additional results in Appendix C. Section C.1 reports training durations for self-supervised pretraining, BiSSL and fine-tuning. Section C.3 presents ablations on BiSSL by lowering the number of upper-level iterations to $N_{U}=1$, underlining the benefits of increasing this value in practice. A sensitivity analysis of $\lambda$ is provided in Section C.4, which reports how classification accuracy is affected by varying this parameter. Section C.5 presents ImageNet fine-tuning results, and Section C.6 demonstrates that BiSSL also achieves significant performance improvements using the official SimCLR-pretrained weights from Chen et al. (2020) with the original $224\crossproduct 244$ image resolution.
5 Exploratory Backbone Alignment Analyses
To assess the extent of which BiSSL improves backbone alignment for fine-tuning, we compare downstream alignment of self-supervised with BiSSL-trained backbones using quantitative and qualitative approaches.
5.1 Measuring Backbone Similarity
We first aim to quantitatively measure the backbone alignment. This is approached by calculating the cosine similarity (CS), firstly between backbone parameters and secondly between backbone output features. Specifically, we compare the CS between SimCLR-pretrained backbones and their fine-tuned versions, with the CS of BiSSL-trained backbones and their fine-tuned counterparts. CS values are computed for models trained on each of the 12 datasets listed in Table 1, with the average CS across these datasets reported in Figure 3. Further experimental setup details are provided in Section B.10.1, where we also provide individual dataset CS values in Table 9. The results imply that BiSSL backbones are consistently more aligned with their fine-tuned versions than standard SSL-pretrained backbones, implying that BiSSL-trained backbones impose a smaller parameter and feature shift during fine-tuning.
@@ 5.2 Visual Inspection of Latent Features
To better assess whether BiSSL nudges the latent features toward being more semantically meaningful for downstream tasks, we qualitatively inspect latent spaces using the t-Distributed Stochastic Neighbor Embedding (t-SNE) (Cieslak et al., 2020) method. Specifically, we compare features from SimCLR-pretrained backbones to those derived from BiSSL-trained lower-level backbones. Further implementation details are provided in Section B.10.2. Figure 4 illustrates the results on the flowers dataset, indicating that BiSSL improves downstream feature alignment. Additional plots on a selection of downstream datasets in Section C.7 reinforce this finding, demonstrating that this trend consistently persists.
Figure 4: Visualizations of backbone output features on the Flowers dataset for the self-supervised pretrained backbone (left) and the BiSSL-trained backbone (right). Colors indicate different classes. Further details are provided in Section B.10.2, and analogous visualizations for other datasets are found in Section C.7. The BiSSL backbone output features exhibit improved discriminative alignment with the downstream classes.
@@ 6 Conclusion
This study introduces BiSSL, a novel training framework that integrates the self-supervised pretext and downstream fine-tuning objectives into a unified bilevel optimization problem. The bilevel structure enables the downstream objective to guide the pretext task optimization towards refining its representations to better align with the downstream task. We present a task-agnostic training algorithm and pipeline that incorporates BiSSL as a preparatory stage prior to fine-tuning. Experiments across multiple pretext and downstream tasks demonstrate that BiSSL consistently enhances fine-tuning performance, while additional alignment analyses indicate that it produces pretrained representations that better conform to downstream objectives. Overall, BiSSL demonstrates that explicitly addressing the misalignment between self-supervised pretrained backbones and downstream tasks prior to fine-tuning can yield substantial accuracy gains. This highlights the potential of bilevel optimization in this setting and motivates future research on training algorithms that explicitly optimize such alignment.
Broader Impact Statement
Our work focuses on improving downstream performance of self-supervised pretrained models without introducing new societal or ethical risks beyond those already associated with standard self-supervised learning and fine-tuning.
Acknowledgments
@@
We acknowledge the Danish e-Infrastructure Consortium (DeiC) for awarding this project access to the LUMI supercomputer. Lastly, the authors would like to thank Sijia Liu and Yihua Zhang (Michigan State University) for providing valuable feedback in a discussion, which helped to refine and solidify our perspective on the topic of integrating bilevel optimization in deep learning.
@@ Appendix A Theoretical Insights and Framework Comparisons in BiSSL
A.1 Derivation of the Implicit Jacobian
@@
$G({\bf\it\theta}_{D},{\bf\it\theta}_{P}):=\mathcal{L}^{P}({\bf\it\theta}_{P},{\bf\it\phi}_{P})+\lambda r({\bf\it\theta}_{P},{\bf\it\theta}_{D}).$ (8)
Now, furthermore assume that the regularization objective $r$ is strongly convex, and that $\mathcal{L}^{P}$ is twice-differentiable and $L$-smooth (i.e. it has Lipschitz-continuous gradients). As established and similarly exploited in prior works (Rajeswaran et al., 2019; Zhang et al., 2022b; 2023a), these properties imply that an adequate scaling of $\lambda$ renders the lower-level objective $G$ effectively convex. This is advantageous because assuming convexity of $G$ ensures that for any ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$, there exists a corresponding $\hat{{\bf\it\theta}}_{P}\in\mathbb{R}^{L}$ that satisfies the stationary condition $\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},\hat{{\bf\it\theta}}_{P})=\mathbf{0}$. In other words, we are assured that a potential minimizer of $G({\bf\it\theta}_{D},\cdot)$ exists for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. Now, further assume that the Hessian matrix $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},\hat{{\bf\it\theta}}_{P})$ is invertible for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. Under these conditions, the implicit function theorem (Dontchev and Rockafellar, 2014; Zucchet and Sacramento, 2022) guarantees the existence of an implicit unique and differentiable function ${\bf\it\theta}_{P}^{\ast}:\mathcal{N}({\bf\it\theta}_{D})\to\mathbb{R}^{L}$, with $\mathcal{N}({\bf\it\theta}_{D})$ being a neighborhood of ${\bf\it\theta}_{D}$, that satisfies ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})=\hat{{\bf\it\theta}}_{P}$ and $\nabla_{{\bf\it\theta}_{P}}G(\tilde{{\bf\it\theta}}_{D},{\bf\it\theta}_{P}^{\ast}(\tilde{{\bf\it\theta}}_{D}))=\mathbf{0}$ for all $\tilde{{\bf\it\theta}}_{D}\in\mathcal{N}({\bf\it\theta}_{D})$.
@@
is well-posed for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$. By applying the chain rule, the expression becomes
@@
Recalling that $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))$ is assumed invertible, the implicit Jacobian (IJ) $\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ can be isolated
@@
$\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ $=-\lambda\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\theta}_{D})\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}\quantity(\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+\lambda r\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\theta}_{D}))]^{-1}$
$=-\nabla_{{\bf\it\theta}_{D}{\bf\it\theta}_{P}}^{2}r({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\theta}_{D})\quantity[\nabla_{{\bf\it\theta}_{P}}^{2}\quantity(\frac{1}{\lambda}\mathcal{L}^{P}\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})+r\quantity({\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D}),{\bf\it\theta}_{D}))]^{-1}.$ (9)